# Mining experimental data from Materials Science literature with Large Language Models

### v1

@@ Abstract
This study is dedicated to evaluating the capabilities of advanced large language models (LLMs) such as GPT-3.5-Turbo, GPT-4, and GPT-4-Turbo in the extraction of structured information from scientific documents within the field of materials science. We introduce a novel methodology for the comparative analysis of intricate material expressions, emphasising the standardisation of chemical formulas to tackle the complexities inherent in materials science information assessment. To this end, we primarily focus on two critical tasks of information extraction: (i) a named entity recognition (NER) of studied materials and physical properties and (ii) a relation extraction (RE) between these entities. The performance of LLMs in executing these tasks is benchmarked against traditional models based on the BERT architecture and rule-based approaches. For NER, LLMs fail to outperform the baseline with zero-shot prompting and exhibit only limited improvement with few-shot prompting. However, for RE, a GPT-3.5-Turbo fine-tuned with the appropriate strategy outperforms all models, including the baseline. Without any fine-tuning, GPT-4 and GPT-4-Turbo display remarkable reasoning and relationship extraction capabilities after being provided with merely a couple of examples, surpassing the baseline. Overall, the results suggest that although LLMs demonstrate relevant reasoning skills in connecting concepts, for tasks requiring extracting complex domain-specific entities like materials, specialised models are currently a better choice.
@@ 1 Introduction
Mining experimental data from literature has become increasingly popular in materials science due to the vast amount of information available and the need for accelerating materials discovery using data-driven techniques. Data for machine learning in materials science is often sourced from published papers, material databases, laboratory experiments, or first-principles calculations [1]. The introduction of big data in materials research has led to a shift from traditional random techniques to more efficient, data-driven methods. Data mining of computational screening libraries has been shown to identify different classes of strong $\rm CO_{2}$-binding sites, enabling materials to exhibit specific properties even in wet flue gases [2]. Machine learning techniques have been employed for high-entropy alloy discovery, with a focus on probabilistic models and artificial neural networks [3]. However, the use of advanced machine learning algorithms in experimental materials science is limited by the lack of sufficiently large and diverse datasets amenable to data mining [4]. A present central tenet of data-driven materials discovery is that with a sufficiently large volume of accumulated data, and suitable data-driven techniques, the process of designing a new material could be more efficient and rational [5]. The materials science field is indeed moving away from traditional manual, serial, and human-intensive work towards automated, parallel, and iterative processes driven by artificial intelligence, simulation, and experimental automation [6, 7]. But, materials science literature is a vast source of knowledge that remains relatively unexplored with data mining techniques [8], especially for the reason that materials science data come in diverse forms such as unstructured textual content and structured tables and graphs, adding complexity to the extraction process. As a result, many projects still depend nowadays on manual data extraction. While there exist extensive structured databases containing accumulated experimental data [9], they remain limited in number and extremely costly due to the amount of human labour involved [10].
@@
Additionally, addressing issues related to the quality and meaning of materials science data often demands a curation step assisted by a sub-domain knowledge often specific to the approached sub-field of materials science, e.g. polymers, metal-organic frameworks, high-entropy alloys, etc, with their own physical and chemical phenomena, methods and protocols, terminology and jargon. For instance, the classification of superconductors can be complex and sometimes arbitrary, blending compound-based classes like cuprates [11] and iron-based [12] materials with unconventional classes like heavy fermions [13]. The classification of superconductors can also be based on phenomena such as the Meissner effect, which describes how superconductors expel magnetic fields [14]. Superconductors can be divided into two classes according to how this breakdown occurs, the so-called type-I and type-II superconductors. As these classifications are not mutually exclusive certain materials could potentially fall into multiple categories, for example, a material can be both a cuprate and a type-II superconductor, the classification of superconductors is a complex task that demands an extensive knowledge of the developments and current state-of-the-arts. Moreover, substantial confusion may occur due to the cross-domain polysemy of used words, terms and symbols. In different sub-domains, the same term can take on specific nuances or meanings unique to a given sub-domain. This phenomenon is common in language and can lead to misunderstandings if the context of the sub-domain is not clear. For instance, the acronym ”TC” or ”Tc” will be employed for denoting a ”Temperature Curie” or a ”superconducting critical temperature”, respectively. These sub-domain-specific conventions pose a significant challenge when attempting to effectively create structured datasets across various sub-domains.
@@
Meanwhile, the advent of large language models (LLMs) has inaugurated a new technological era, marked by extraordinary potential. These models not only excel in linking diverse concepts but also in engaging in sophisticated conversational reasoning [15, 16, 17, 18]. Compared to Small Language Models (SLMs), LLMs offer the possibility of integrating large corpus of textual data at training, with often the ability to ingest large textual inputs at inference with a context window ranging from 4,096 to 128,000 tokens for GPT-3.5-Turbo and GPT-4-Turbo [19], respectively (at the time of this writing). By comparison, BERT-based encoders are limited to only 512 tokens, where 1,000 tokens are about 750 English words. Furthermore, the interaction with LLMs via prompts, i.e., tailored instructions, changes the construction paradigm of programmatic solutions which becomes more accessible, flexible and suitable to human operators. Nevertheless, the actual capabilities of LLMs in reasoning, understanding and recognition are still under constant evolution and evaluation.
@@
Previous studies in Information Extraction (IE) have shown evidence of LLMs proficiency in general tasks, presenting a valuable opportunity to develop more flexible Text and Data Mining (TDM) processes, but they fall short in areas where specific knowledge is required [20]. In particular, LLMs are on par with SLMs in most of the discriminative tasks such as named entity recognition (NER), relation extraction (RE) and event detection (ED) in general domains [21], in history [22], and biology [23]. Other works testing chemistry capabilities found that GPT-4 understands various aspects of chemistry, including chemical compounds [24]; however, its knowledge is general and lacks methods for learning through retrieving recent literature [25].
@@
Therefore, this study aims to assess LLMs’ ability to comprehend, manipulate, and reason with complex information that demands substantial background knowledge, as in materials science.
@@
We first classify the fundamental components of the materials science knowledge directed towards the design of novel materials with functional properties into two main entity classes: material and property expressions. Properties, e.g., a critical temperature of 4K, are expressed using measurements of physical quantities. They exhibit a structured format, including modifiers (e.g., ”between”, ”less than”, ”approximately”, or symbols such as ”¿” or ” ”), values, and units, with a wide range of potential values. In contrast, material definitions are conceptually loose and often depend on the specific domain. They may require a substantial amount of accompanying text for a comprehensive description, encompassing details, e.g., compositional ratios, doping agent and ratio, synthesis protocol, process, and additional adjunct information. From a fundamental compositional standpoint, materials are defined by their chemical formula. However, in practice, authors in literature may frequently employ substantives such as commercial names, well-known terms, or crafted designations to describe samples, all of which streamline information in their research papers. Nonetheless, conveying such definitions can unambiguously be challenging.
@@
We address Q2 by assessing the capability to establish connections between a predefined set of entities and extract relationships within a given context.
@@
The evaluation of generative models brings an additional complexity. Traditional SLMs implementations for solving NER tasks are based on sequence labelling algorithms. They classify each token in the input stream with a limited number of labels, returning a sequence that fits perfectly the original input (same number of tokens and structure). Evaluating their performance against expected datasets involves a straightforward comparison of values. Soft-matching techniques can be employed to overlook minor discrepancies. However, with generative models, the output tokens may be structured in ways that significantly differ from the original input sequence. In more general scenarios, semantic models that compare the vectorised representations of two sequences can be utilised [28]. Nevertheless, when dealing with concepts like material expressions, a specialised approach is needed. As an illustration, the terms ”solar cell” and ”solar cells” represent identical concepts, yet the materials denoted by ”Ca” (Calcium) and ”Cr” (Chromium) are entirely distinct, highlighting a difference of just one letter between the two examples. For this reason, we introduce a novel evaluation method for material names, which involves normalising materials to their chemical formulas before conducting a pairwise comparison of each element. This approach provides a more meaningful and context-aware assessment of the model’s performance.
@@
  • We designed and ran a benchmark for LLMs on information extraction, in particular NER of materials and properties. This contribution addresses Q1.
@@ 2 Method
We chose three OpenAI LLM models reported with their specific names for performing API calls: GPT-3.5-Turbo (gpt-3.5-turbo-0611), GPT-4 (gpt-4), and GPT-4-Turbo (gpt-4-0611-preview). The consideration of open-source LLMs has been deferred to future work due to their present limited capability to generate output in a valid JSON format (Section 2.1.1, necessitating a more in-depth investigation.
Our evaluation is carried out with different strategies: zero-shot prompting, few-shot prompting and fine-tuning (or instruction-learning). Few-shot prompting refers to the model’s ability to adapt and perform a new task with minimal examples or prompts, while zero-shot prompting denotes the model’s capability to generalise to tasks it has not been explicitly trained on, emphasising transfer learning within the language domain. Finally, fine-tuning involves adjusting the parameters of a pre-trained model on a specific task or domain using a smaller, task-specific dataset to enhance its performance for that particular application.
We selected two datasets for evaluation: MeasEval [29], a SemEval 2021 task of extracting counts, measurements, and related context from scientific documents and SuperMat, an annotated and linked dataset of research papers on superconductors [30]. SuperMat contains both materials and properties and for copyright reasons is not publicly distributed. This reduces the risk that its annotations had been used during the pre-training of any of the LLMs.
@@
The presented evaluations condense average F1 scores and their standard deviation over three extraction runs. Detailed values are available in the Supplementary Materials section.
@@ 2.1 Named Entities Recognition
We calculated the evaluation scores using four different matching approaches, although we will present only the most relevant to the task (leaving the complete tables [^1] The calculation of micro average provides a measure independent of the distribution of the extracted entities over the different documents. in the supplementary materials section):
@@
  • soft Matching using Ratcliff/Obershelp [^2] https://www.drdobbs.com/database/pattern-matching-the-gestalt-approach/184407970?pgno=5 with a threshold at 0.9333The threshold is fixed to a value yielding more than 90% similarity.
  • Sentence BERT Comparison using semantic similarity of sequences using Sentence BERT with a cross-encoder [28], applying a threshold set at 0.9\@textsuperscript3
@@
Prompts for interacting with LLMs are defined by two components: system and user prompts. The system prompt is the initial instruction guiding the model’s output generation, defining the task or information sought, while the user prompt is the input from the user, specifying their request and shaping the model’s response.
The system prompt below was fixed across all tasks and was specifically crafted to prevent the making of non-existing facts and favour standardised answers (e.g., ”I don’t know”, ”None”) in case of inability to respond.
@@
Listing 2: User prompt designed for extracting materials and properties. They are separated by ”——”.
@@
Include relevant text that indicates the application of a modifier, such as "between" "less than" "approximately",
@@
Then, we applied a few-shot prompting technique by incorporating in the users’ prompt template above a set of suggestions extracted from the text (see Listing 3 below) using the respective SLMs based on the fine-tuned SciBERT-encoder for materials and properties, i.e., grobid-superconductors [26] and grobid-quantities [27], respectively. Also, as these suggestions originate from another model, they may not be entirely accurate; hence, we emphasised in the prompts that they only serve as examples or hints that may be ignored by the LLMs.
@@ 2.1.1 Output format
For all tasks, we required the output to be formatted using a valid JSON document. We justify this decision for three main reasons: a) The responses need to be machine-readable so that the de-serialisation from JSON to objects in many programming languages becomes a trivial operation (e.g., Python, JavaScript). b) The JSON schema can be defined through a documented format regardless of the programming language or platform. Finally, c) the JSON format, being an open standard, can be used by anyone and does not require reinventing the wheel by re-implementing from scratch any of the transformation steps.
The JSON output was obtained by adding formatting instructions in the user’s prompt based on the expected output data model, for which different concepts were described in different ways (e.g., properties are described as a value and an optional unit). We used the implementation provided by the LangChain library [^6] https://github.com/langchain-ai/langchain of which one example is illustrated in the following listing.

### v3

@@ Abstract
This study is dedicated to assessing the capabilities of large language models (LLMs) such as GPT-3.5-Turbo, GPT-4, and GPT-4-Turbo in extracting structured information from scientific documents in materials science. To this end, we primarily focus on two critical tasks of information extraction: (i) a named entity recognition (NER) of studied materials and physical properties and (ii) a relation extraction (RE) between these entities. Due to the evident lack of datasets within Materials Informatics (MI), we evaluated using SuperMat, based on superconductor research, and MeasEval, a generic measurement evaluation corpus. The performance of LLMs in executing these tasks is benchmarked against traditional models based on the BERT architecture and rule-based approaches (baseline). We introduce a novel methodology for the comparative analysis of intricate material expressions, emphasising the standardisation of chemical formulas to tackle the complexities inherent in materials science information assessment. For NER, LLMs fail to outperform the baseline with zero-shot prompting and exhibit only limited improvement with few-shot prompting. However, a GPT-3.5-Turbo fine-tuned with the appropriate strategy for RE outperforms all models, including the baseline. Without any fine-tuning, GPT-4 and GPT-4-Turbo display remarkable reasoning and relationship extraction capabilities after being provided with merely a couple of examples, surpassing the baseline. Overall, the results suggest that although LLMs demonstrate relevant reasoning skills in connecting concepts, specialised models are currently a better choice for tasks requiring extracting complex domain-specific entities like materials. These insights provide initial guidance applicable to other materials science sub-domains in future work.
@@ 1 Introduction
Mining experimental data from literature has become increasingly popular in materials science due to the vast amount of information available and the need to accelerate materials discovery using data-driven techniques. Data for machine learning in materials science is often sourced from published papers, material databases, laboratory experiments, or first-principles calculations [1]. The introduction of big data in materials research has shifted from traditional random techniques to more efficient, data-driven methods. Data mining of computational screening libraries has been shown to identify different classes of strong $\rm CO_{2}$-binding sites, enabling materials to exhibit specific properties even in wet flue gases [2]. Machine learning techniques have been employed for high-entropy alloy discovery, focusing on probabilistic models and artificial neural networks [3]. However, the use of advanced machine learning algorithms in experimental materials science is limited by the lack of sufficiently large and diverse datasets amenable to data mining [4]. A present central tenet of data-driven materials discovery is that with a sufficiently large volume of accumulated data and suitable data-driven techniques, designing a new material could be more efficient and rational [5]. The materials science field is moving away from traditional manual, serial, and human-intensive work towards automated, parallel, and iterative processes driven by artificial intelligence, simulation, and experimental automation [6, 7]. But, materials science literature is a vast source of knowledge that remains relatively unexplored with data mining techniques [8], especially for the reason that materials science data come in diverse forms such as unstructured textual content and structured tables and graphs, adding complexity to the extraction process. As a result, nowadays, many projects still depend on manual data extraction. While extensive structured databases contain accumulated experimental data [9], they remain limited in number and highly costly due to the amount of human labour involved [10].
@@
Additionally, addressing issues related to the quality and meaning of materials science data often demands a curation step assisted by a sub-domain knowledge frequently specific to the approached sub-field of materials science, e.g. polymers, metal-organic frameworks, high-entropy alloys, etc., with their own physical and chemical phenomena, methods and protocols, terminology and jargon. For instance, the classification of superconductors can be complex and sometimes arbitrary, blending compound-based classes like cuprates [11] and iron-based [12] materials with unconventional classes like heavy fermions [13]. The classification of superconductors can also be based on phenomena such as the Meissner effect, which describes how superconductors expel magnetic fields [14]. Superconductors can be divided into two classes according to how this breakdown occurs, the so-called type-I and type-II superconductors. As these classifications are not mutually exclusive certain materials could potentially fall into multiple categories, for example, a material can be both a cuprate and a type-II superconductor, the classification of superconductors is a complex task that demands an extensive knowledge of the developments and current state-of-the-arts. Moreover, substantial confusion may occur due to the cross-domain polysemy of used words, terms and symbols. In different sub-domains, the same term can take on specific nuances or meanings unique to a given sub-domain. This phenomenon is common in language and can lead to misunderstandings if the context of the sub-domain is unclear. For instance, the acronym ”TC” or ”Tc” will be employed for denoting a ”Temperature Curie” or a ”superconducting critical temperature”, respectively. These sub-domain-specific conventions pose a significant challenge when attempting to create structured datasets across various sub-domains effectively.
Meanwhile, the advent of large language models (LLMs) has inaugurated a new technological era marked by extraordinary potential. These models not only excel in linking diverse concepts but also in engaging in sophisticated conversational reasoning [15, 16, 17, 18]. In comparison, rule-based approaches are simpler and faster (tokens per second); however, they are time-intensive to fine-tune and have weak generalisation capabilities, as new rules should be clarified on a case-by-case basis. Small Language Models (SLMs), e.g. BERT-based models, are more specific to the task on which they are pre-trained. The data size used for pre-training LLMs is usually high enough to contain a high diversity of contexts, thus necessitating fewer examples at the fine-tuning stage than SLMs models.
@@
LLMs offer the possibility of integrating large corpus of textual data at training, with often the ability to ingest large textual inputs at inference with a context window ranging from 4,096 to 128,000 tokens for GPT-3.5-Turbo and GPT-4-Turbo [19], respectively (at the time of this writing). Differently, BERT-based encoders are limited to only 512 tokens, whereas 1,000 tokens are about 750 English words. The size of BERT models does not allow them to sustain their contextual memory after fine-tuning. LLMs possess the capacity to be fine-tuned and retain the contextual knowledge from pre-training, which gives them an advantage in terms of generalisation to other datasets. Finally, the interaction with LLMs via prompts, i.e., tailored instructions, changes the construction paradigm of programmatic solutions, making them more accessible, flexible, and suitable to human operators. Nevertheless, the actual capabilities of LLMs in reasoning, understanding, and recognition are still constantly evolving and being evaluated.
@@
Previous studies in Information Extraction (IE) have shown evidence of LLMs proficiency in general tasks, presenting a valuable opportunity to develop more flexible Text and Data Mining (TDM) processes. Still, they fall short in areas where specific knowledge is required [20]. In particular, LLMs are on par with SLMs in most of the discriminative tasks such as named entity recognition (NER), relation extraction (RE) and event detection (ED) in general domains [21], in history [22], and biology [23]. Other works testing chemistry capabilities found that GPT-4 understands various aspects of chemistry, including chemical compounds [24]; however, its knowledge is general and lacks methods for learning through retrieving recent literature [25].
@@
Therefore, this study assesses LLMs’ ability to comprehend, manipulate, and reason with complex information that demands substantial background knowledge, as in materials science.
@@
We first classify the fundamental components of the materials science knowledge directed towards designing novel materials with functional properties into two main entity classes: material and property expressions. Properties, e.g., a critical temperature of 4 K, are expressed using measurements of physical quantities. They exhibit a structured format, including modifiers (e.g., ”between”, ”less than”, ”approximately”, or symbols such as ”$>$” or $\sim$”), values, and units, with a wide range of potential values. In contrast, material definitions are conceptually loose and often depend on the specific domain. They may require a substantial amount of accompanying text for a comprehensive description, encompassing details, e.g., compositional ratios, doping agent and ratio, synthesis protocol, process, and additional adjunct information. From a fundamental compositional standpoint, materials are defined by their chemical formula. However, in practice, authors in literature may frequently employ substantives such as commercial names, well-known terms, or crafted designations to describe samples, all of which streamline information in their research papers. Nonetheless, conveying such definitions can unambiguously be challenging.
@@
Named Entity Recognition (NER) [26], alternatively referred to as named entity identification or entity extraction, stands as a pivotal component within information extraction. Its primary objective is to pinpoint and categorise named entities within unstructured text, assigning them to predefined categories such as individual names, organisations, geographic locations, medical codes, temporal expressions, quantities, monetary values, percentages, and more. The process of identifying entities aligns closely with sequence labelling tasks, wherein a string of text undergoes analysis, and each token within it (basic unit of text processing, typically a word or a sequence of characters that is treated as a single unit) is designated to one of the pre-established categories. For instance, these categories may include material, doping, condition, or property, among others.
We address Q2 by assessing the capability to establish connections between a predefined set of entities and extract relationships within a given context. Extracting relations between entities is a foundational undertaking in NLP. It entails discerning connections or associations among entities referenced within textual data. For instance, in biomedical research, relationship extraction might involve identifying the association between specific genes and diseases mentioned in scientific literature.
@@
The evaluation of generative models brings an additional complexity. Traditional SLM implementations for solving NER tasks are based on sequence labelling algorithms. They classify each token in the input stream with a limited number of labels, returning a sequence that fits the original input (same number of tokens and structure). Evaluating their performance against expected datasets involves a straightforward comparison of values. Soft-matching techniques can be employed to overlook minor discrepancies. However, with generative models, the output tokens may be structured in ways that significantly differ from the original input sequence. In more general scenarios, semantic models that compare the vectorised representations of two sequences can be utilised [29]. Nevertheless, when dealing with concepts like material expressions, a specialised approach is needed. As an illustration, the terms ”solar cell” and ”solar cells” represent identical concepts. Yet, the materials denoted by ”Ca” (Calcium) and ”Cr” (Chromium) are entirely distinct, highlighting a difference of just one letter between the two examples. For this reason, we introduce a novel evaluation method for material names, which involves normalising materials to their chemical formulas before conducting a pairwise comparison of each element. This approach provides a more meaningful and context-aware assessment of the model’s performance.
@@
  • We designed and ran a benchmark for LLMs on information extraction, particularly NER of materials and properties. This contribution addresses Q1.
@@ 2 Method
We chose three OpenAI LLM models reported with their specific names for performing API calls: GPT-3.5-Turbo (gpt-3.5-turbo-0611), GPT-4 (gpt-4), and GPT-4-Turbo (gpt-4-0611-preview). The consideration of open-source LLMs has been deferred to future work due to their limited capability to generate output in a valid JSON format (Section 2.1.1, necessitating a more in-depth investigation.
Our evaluation uses different strategies: zero-shot prompting, few-shot prompting, and fine-tuning (or instruction-learning). Few-shot prompting refers to the model’s ability to adapt and perform a new task with minimal examples or prompts. In contrast, zero-shot prompting denotes the model’s capability to generalise to tasks it has not been explicitly trained on, emphasising transfer learning within the language domain. Finally, fine-tuning involves adjusting the parameters of a pre-trained model on a specific task or domain using a smaller, task-specific dataset to enhance its performance for that particular application.
We selected two datasets for evaluation: MeasEval [30], a SemEval 2021 task of extracting counts, measurements, and related context from scientific documents and SuperMat, an annotated and linked dataset of research papers on superconductors [31]. SuperMat contains both materials and properties and, for copyright reasons, is not publicly distributed. This reduces the risk that its annotations had been used during the pre-training of any of the LLMs.
@@
The evaluations condense average F1 scores and their standard deviation over three extraction runs. The raw tables with all detailed scores are provided in Appendix A.
@@ 2.1 Named Entities Recognition
We calculated the evaluation scores using four different matching approaches. However, we will present only the most relevant to the task (leaving the complete tables [^1] The calculation of micro average provides a measure independent of the distribution of the extracted entities over the different documents. in Appendix A):
@@
  • soft Matching using Ratcliff/Obershelp [33] with a threshold at 0.9222The threshold is fixed to a value yielding more than 90% similarity.
  • Sentence BERT Comparison using semantic similarity of sequences using Sentence BERT with a cross-encoder [29], applying a threshold set at 0.92
@@
Prompts for interacting with LLMs are defined by two components: system and user prompts. The system prompt is the initial instruction guiding the model’s output generation, defining the task or information sought. In contrast, the user prompt is the user’s input, specifying their request and shaping the model’s response.
The system prompt below was fixed across all tasks. It was specifically crafted to prevent the creation of non-existing facts and favour standardised answers (e.g., ”I don’t know,” ”None,” etc.) in case of inability to respond.
@@
Listing 2: User prompt designed for extracting materials and properties. The entity descriptions are separated by dashed lines (”------”).
@@
Include relevant text that indicates the application of a modifier, such as "between", "less than", "approximately",
@@
Then, we applied a few-shot prompting technique by incorporating in the users’ prompt template above a set of suggestions extracted from the text (see Listing 3 below) using the respective SLMs based on the fine-tuned SciBERT-encoder for materials and properties, i.e., grobid-superconductors [27] and grobid-quantities [28], respectively. Also, as these suggestions originate from another model, they may not be entirely accurate; hence, we emphasised in the prompts that they only serve as examples or hints that the LLMs may ignore.
@@ 2.1.1 Output format
For all tasks, we required the output to be formatted using a valid JSON document. We justify this decision for three main reasons: a) The responses need to be machine-readable so that the de-serialisation from JSON to objects in many programming languages becomes a trivial operation (e.g., Python, JavaScript). b) The JSON schema can be defined through a documented format regardless of the programming language or platform. Finally, c) the JSON format is an open standard that can be used by anyone and does not require reinventing the wheel by re-implementing any transformation steps from scratch.
The JSON output was obtained by adding formatting instructions in the user’s prompt based on the expected output data model, for which different concepts were described differently (e.g., properties are described as a value and an optional unit). We used the implementation provided by the LangChain library [^5] https://github.com/langchain-ai/langchain of which one example is illustrated as follows.